OpenStudy (anonymous):

Last question. "Show (by expanding) that sqrt(2) * sin[x + (pi / 4)] = sinx + cosx."

OpenStudy (anonymous):

I'm not really sure where to start. But I've expanded \[\sin (x + \frac{ \pi }{ 4 })\] using the two-angle formula.

OpenStudy (raden):

yes that's right, just continue it.. do you know what's the value of sin(pi/4) and cos(pi/4) ?

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{2} }\]

OpenStudy (raden):

correct! then simplify ....

OpenStudy (anonymous):

Oh, I understand now! I see how the sqrt(2) cancels with those now.

OpenStudy (anonymous):

Thank you for that! I'll give it another go.

OpenStudy (raden):

you're welcome